OpenStudy (anonymous):

Calc 2 the half life of C-14 is about 5730 years a. archaeologists find a piece of cloth painted with organic dyes. Analysis of the dye in the cloth shows that only 77% of the c-14 originally in the dye remains. When was the cloth painted? b. a well preserved piece of wood found at an archaeological site has 6.2% of the C-14 that it had when it was alive. Estimate when the wood was cut.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

don't need any calc for this one

OpenStudy (anonymous):

solve \[\large \left(\frac{1}{2}\right)^{\frac{t}{5700}}=.77\] for \(t\)

OpenStudy (anonymous):

and b?

OpenStudy (anonymous):

you know how to solve the first one ?

OpenStudy (anonymous):

oh damn i didn't read carefully it is \[\large \left(\frac{1}{2}\right)^{\frac{t}{5730}}=.77\]

OpenStudy (anonymous):

second is identical to the first, but replace \(.77\) by \(.062\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw